OpenStudy (anonymous):

Find the coordinates for the center and length of radius for this circle. (x + 3)2 + (y - 5)2 = 25 center ( ,) radius ____

Parth (parthkohli):

In an equation like this: \(\Large \color{Black}{\Rightarrow (x - h)^2 + (y - b)^2 = r^2 }\) r is the radius. (h,b) are the center co-ordinates.

Parth (parthkohli):

(-3,5) are center co-ordinates and 5 is the radius.

OpenStudy (anonymous):

thanks!

Parth (parthkohli):

5 is the radius because r^2 = 25 and r = 5

Parth (parthkohli):

Yw :)

OpenStudy (anonymous):

wait isnt it (3,-5)?

Parth (parthkohli):

Nope, it would've been see if the equation of a circle was (x + h)^2 + (y + b)^2 = r^2

OpenStudy (anonymous):

oh ok:D

Parth (parthkohli):

Here it is (x - h)^2 + (y - b)^2

Parth (parthkohli):

:P

OpenStudy (anonymous):

(x - 3)2 + (y - 5)2 = 8 how about here^^ now would it be -3 ,-5??

Parth (parthkohli):

No, 3,5 only

OpenStudy (anonymous):

or both positive 3,5?

OpenStudy (anonymous):

oh ok:)

Parth (parthkohli):

Yep...positive